### v2

# Topology of dyonic AdS black holes with quasitopological electromagnetism in Einstein-Gauss-Bonnet gravity

Abstract
Abstract: In this study, we employ the thermodynamic topological method to classify critical points for the dyonic AdS black holes with quasitopological electromagnetism in the Einstein-Gauss-Bonnet background. To this end, we find a small/large black hole phase transition in all dimensions of space-time, the existence of a conventional critical point implies a total topological charge of $Q_{t}=-1$. The coupling constant $\alpha$ gives rise to a more intricate phase structure, with the emergence of a triple points requires $\alpha\geq 0.5$ and $d=6$. Interestingly, the condition for the occurrence of small/intermediate/large phase transition is that the coupling constant a takes a special value ($\alpha=0.5$), the two conventional critical points $CP_{1},CP_{2}$ of the black hole are physical critical point, and the novel critical point $CP_{3}$ that lacks the capability to minimize the Gibbs free energy. The critical points $CP_{1}$ and $CP_{2}$ are observed to occur at the maximum extreme points of temperature in the isobaric curve, while the critical point $CP_{3}$, emerges at the minimum extreme points of temperature. Furthermore, the number of phases at the novel critical point exhibits an upward trend, followed by a subsequent decline at the conventional critical points. With the increase of the coupling constant ($\alpha=1$), although the system has three critical points, only the conventional $CP_{1}$ is a (physical) critical point, and the conventional $CP_{2}$ serves as the phase annihilation point. This means that the coupling constant $\alpha$ has a non-negligible effect on the phase structure.
preprint: APS/-QNM
I Introduction
The official announcement on February 11, 2016 of the detection of gravitational waves by the LIGO represents a groundbreaking achievement that provides significant validation for Einstein’s general relativity. This momentous discovery not only confirms the existence of black holes but also verifies the presence of gravitational waves ch1 ; ch2 ; ch3 ; ch4 . The emergence of black hole (BH) thermodynamics stems from the application of general laws in thermodynamics to the general relativity, with the objective of investigating the characteristics inherent to black holes. The fundamental concept involves regarding black holes as a system governed by thermodynamic principles, wherein mass corresponds to energy, entropy is directly proportional to the surface area, and temperature is represented by surface gravity. In 1974, Hawking made a groundbreaking discovery that black holes exhibit thermal radiation ch5 , this revelation not only resolves the inherent contradictions within the thermodynamics of black holes but also successfully establishes a profound connection between gravitational theory, quantum mechanics, and thermodynamics. The phenomenon of phase transition is commonly observed in thermodynamic systems, wherein the Schwarzschild black hole, when considered as a thermodynamic system within the framework of AdS, exhibits an unstable state from a thermodynamic perspective at low temperatures. However, with increasing temperature, the free energy is reduced, and finally the black hole reaches stability. Essentially, the stability of a black hole hinges on the critical temperature known as the Hawking-Page (HP) phase transition ch6 . Based on the AdS/CFT correspondence theory, there has been considerable focus on studying the HP phase transition, which is recognized as a phase transition between confinement and deconfinement in the gauge/gravity duality ch7 .
Kastor and Ray proposed the concept of regarding the cosmological constant $\Lambda$ as a variable thermodynamic pressure, and suggested that the volume of the black hole in AdS space can be considered as its conjugate quantity, while the black hole’s mass is interpreted as the spacetime’s enthalpy, rather than its internal energy ch8 . This notion was further expanded to Lovelock theory and applied to black holes, which demonstrates that Lovelock coupling serves as an extension of the first law of thermodynamics ch9 . Additionally, both the free energy and of Arnowitt-Deser- Misner (ADM) mass AdS black holes were derived ch10 . Furthermore, investigations on dilaton-AdS black holes have revealed that their mass represents the system’s enthalpy ch11 , the phenomenon has led to a surge in research on thermodynamic phase transitions in AdS black holes. For instance, the charged RN-AdS black hole in four-dimensional spacetime was examined by Kubiznak and Mann, who identified a significant transition between small and large black holes, which bears resemblance to the gas-liquid phase transition observed in van der Waals fluids ch17 . The phase transition in nonlinear electromagnetic field was investigated by Hendi and Vahidinia through considering the Einstein gravitational solution under the background of generalized Maxwell theory ch18 . Critical phenomena and critical exponents near the critical point of Gauss-Bonnet black hole were studied by Wei et al, who also examined the phenomenon of phase transition in spin Kerr black holes with a high number of dimensions ch19 ; ch20 . Zou et al. explored P-V critical behavior of Born-Infeld Anti-deSitter black hole and Gauss-Bonnet AdS black hole ch21 ; ch22 . It is well known that the typical first-order phase transition has only one critical point. However, the existence of multiple critical points is observed in certain categories of black holes, such as, the occurrence of reentrant phase transition ch24 ; ch25 , the presence of triple points in dynamic AdS black holes under quasitopological electromagnetism (QTE) ch26 ; ch27 , and the $\lambda$ phase transition observed in Horava gravity ch28 , among others. A complementary aspect to the investigation of black hole criticality is the concept of thermodynamic topology. The Duan ’s $\varphi$-mapping theory proposed is utilized to explore the criticality of black holes, resulting in distinct characteristics of topological charge exhibited by critical points within the thermodynamic space, these critical points can be classified into conventional and novel types cq4 . Furthermore, Wei et al. employed the thermodynamic topological method to investigate thermodynamic properties of black holes, this approach enables analysis of stability by considering the winding number associated with defect we1 . The thermodynamic topology classification of different black holes are found in detail in we2 ; we2.1 ; we3 ; we4 ; we5 ; we7 ; we9 ; we8 ; we10 ; we12 ; we13 ; we14 ; we15 ; we16 ; we17 ; we17.1 ; we17.2 ; we18 .
The theory of extended gravity, which encompasses Einstein-Gauss-Bonnet (EGB) gravity, provides novel perspectives on the characteristics of black holes ch29 ; ch30 ; ch31 ; ch32 ; ch33 ; ch34 ; ch35 . Through an examination of the thermodynamic characteristics of the Born-Infeld AdS black hole within the framework of 4D EGB gravity, it has been observed that there exist three critical points within the black hole, these critical points exhibit numerous intriguing properties when subjected to EGB gravity ch25 . Recently, the QTE is characterized by the square norm of the topological 4-form $F\wedge F$, indicating that the system’s energy-momentum tensor corresponds to a perfect fluid with isotropic properties ch36 . Based on this, the thermodynamic phase transition of dyonic AdS black holes with QTE is discussed in detail. By changing the coupling parameters, a triple points phase structure is observed ch27 . Additionally, through an examination of the interplay between QTE and EGB gravity, Sekhmani et al. derived a solution for dyonic AdS black holes with a high number of dimensions. Furthermore, they conducted a comprehensive analysis on how the coupling constant $\alpha$ influences both the energy emission rates and black hole shadow ch37 . On this basis, a comprehensive analysis is conducted on the phase transition of the dyonic AdS black holes with QTE in the EGB background, and it is found that the system parameters will appear triple points in special regions cq3 . In this work, we will use the thermodynamic topological method to investigate the critical points, and further understand the influence of GB coupling constant on the phase structure of dyonic AdS black holes. The structure of the manuscript is as follows: In section II, we give a concise overview of the thermodynamic properties of dyonic AdS black holes with QTE in the EGB background and present a comprehensive explanation of Duan’s $\varphi$-mapping theory. In section III, we will use the thermodynamic topology method to classify the critical points in different space-time dimensions. In section IV, we determine the characteristics of the critical point by analyzing the isobaric curve. Section V contains the remarks and conclusions.
II Dyonic AdS black holes and topology
We first consider the d-dimensional action involving EGB gravity and QTE which can be formulated in a minimal coupling manner ch37
$S_{D}=\frac{1}{16\pi}\int d^{d}x\sqrt{-g}\left(\mathcal{R}-2\Lambda+\alpha\mathcal{G}+\mathcal{L}_{\mathrm{QE}}\right),$ (1)
wherein, $\Lambda=-\frac{(d-1)(d-2)}{2l^{2}}$ represents the cosmological constant within the AdS space, $\alpha$ is the symbol used to denote the GB coupling constant, which possesses dimensions of [length]${}^{2}$. Here, the GB term can be expressed as
$\mathcal{G}=\mathcal{R}^{2}-4\mathcal{R}^{\mu\nu}\mathcal{R}_{\mu\nu}+\mathcal{R}^{\mu\nu\rho\sigma}\mathcal{R}_{\mu\nu\rho\sigma}.$ (2)
The matter field, a fundamental concept in the realm of physics, can be characterized by an electromagnetic Lagrangian with quasi-topological properties and represented as
$\mathcal{L}_{\rm{QE}}=-(\frac{1}{4}\mathcal{F}^{2}+\frac{1}{2p!}\mathcal{H}^{2}+\beta\mathcal{L}_{\rm{int}}),$ (3)
where $p=d-2$, $\mathcal{F}^{2}=\mathcal{F}_{\mu\nu}\mathcal{F}^{\mu\nu}$ and $\mathcal{H}^{2}=\mathcal{H}_{\rho_{1}\ldots\rho_{p}}\mathcal{H}^{\rho_{1}\ldots\rho_{p}}$. Here, $\beta$ means the coupling constant associated with the dimension [length]${}^{2}$. Additionally, the interaction term reads
$\mathcal{L}_{\rm{int}}=\delta_{\gamma_{1}\ldots\gamma_{D}}^{\lambda_{1}\ldots\lambda_{D}}\mathcal{F}_{\lambda_{1}\lambda_{2}}\mathcal{H}_{\lambda_{3}\ldots\lambda_{D}}\mathcal{F}^{\gamma_{1}\gamma_{2}}\mathcal{H}^{\gamma_{3}\ldots\gamma_{D}}.$ (4)
The above action, referring to a specific mathematical calculation or physical process, leads to the emergence of a BH solution, which can be described by a metric
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}d\Omega_{d-2}^{2},$ (5)
[…]
$f(r)=1+\frac{r^{2}}{2\tilde{\alpha}}\left(1-\sqrt{1+\frac{4\tilde{\alpha}m}{r^{d-1}}+\frac{8\tilde{\alpha}\Lambda}{(d-2)(d-1)}-\varrho}\right),$ (6)
with
$\varrho=\frac{2\tilde{\alpha}}{d-3}\left(Q_{m}^{2}+Q_{e}^{2}\xi\right),$ (7)
the hypergeometric function, denoted by $\xi$, can be expressed as
$\xi=\ _{2}F_{1}\left[1,\frac{d-3}{2(d-2)};\frac{7-3d}{4-2d};\frac{-8\beta Q_{m}^{2}\Gamma(d-1)^{2}}{r^{2d-4}}\right],$ (8)
$\widetilde{\alpha}=(d-4)(d-3)\alpha.$ (9)
Here, $Q_{e}$ and $Q_{m}$ are essential parameters that indicate the electrical and magnetic charges of black holes, respectively. These parameters are critical in comprehending the characteristics of black holes. $m$ denotes the mass of the solution within a designated parameter domain, serving as the constant of integration derived from the boundary conditions ch37 . In this case, ADM mass is given by
$M=\frac{(d-2)\mathcal{V}_{d-2}}{16\pi}m,$ (10)
with
$\mathcal{V}_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)},$ (11)
the symbol $\mathcal{V}_{d-2}$ represents the volume of a unit sphere in $(d-2)$ dimensions. The concept of thermodynamic pressure is closely related to the cosmological constant $\Lambda$, specifically, $P=-\frac{\Lambda}{8\pi}$ ch8 . Based on this, the event horizon’s radius can be utilized to ascertain thermodynamic properties of black holes, including Hawking temperature, entropy, volume, and Gibbs free energy
$T$ $=\frac{1}{\left[8(d-2)\pi r_{+}+16\pi r_{+}^{-1}\mathcal{N}_{1}\right]\mathcal{N}_{2}}[(12+12d(d-5)$ (12)
$\left.\left.+32\pi Pr_{+}^{2}+2d(d-5)r_{+}^{-2}\mathcal{N}_{1}-r_{+}^{6-2d}Q_{m}^{2}\right)\mathcal{N}_{2}-r_{+}^{6}Q_{e}^{2}\right],$
with
$\mathcal{N}_{1}$ $=$ $(d-4)(d-3)(d-2)\alpha,$ (13)
$\mathcal{N}_{2}$ $=$ $r_{+}^{2d}+8r_{+}^{4}\beta Q_{m}^{2}\Gamma(d-1)^{2},$ (14)
[…]
$=\frac{\pi^{\frac{d-1}{2}}r_{+}^{d-2}\left[1+2(d-3)(d-2)\alpha r_{+}^{-2}\right]}{2\Gamma\left(\frac{d-1}{2}\right)},$
$V=\left(\frac{\partial H}{\partial P}\right)_{S,Q_{e},Q_{m}}=\frac{2\pi^{\frac{d-1}{2}}r_{+}^{d-1}}{(d-1)\Gamma\left(\frac{d-1}{2}\right)},$ (16)
$G$ $=2(d-2)r_{+}^{2d+2}\mathcal{N}_{3}+2r_{+}^{2d}\mathcal{N}_{1}\mathcal{N}_{3}+\frac{32\pi P\mathcal{N}_{3}r_{+}^{2d+4}}{d-1}$ (17)
$-\frac{\mathcal{N}_{3}}{\left((d-2)r_{+}^{2}+2\mathcal{N}_{1}\right)\mathcal{N}_{2}}\left[2(d-3)(d-2)\alpha+r_{+}^{2}\right]$
×$\times\left[\left(2\mathcal{N}_{4}r_{+}^{2d}-r_{+}^{8}Q_{m}^{2}\right)\mathcal{N}_{2}-r_{+}^{2d+8}Q_{e}^{2}\right]$
$+\frac{r_{+}^{8}\left(Q_{m}^{2}+Q_{e}^{2}\xi_{r_{+}}\right)\mathcal{N}_{3}}{d-3},$
with
$\mathcal{N}_{3}=\frac{\pi^{\frac{d-3}{2}}r_{+}^{-5-d}}{16\Gamma\left(\frac{d-1}{2}\right)},$ (18)
$\mathcal{N}_{4}=6r_{+}^{2}+d(d-5)r_{+}^{2}+16P\pi r_{+}^{4}+(d-5)\mathcal{N}_{1}.$ (19)
To gain a deeper comprehension of the phase structure, we will employ the thermodynamic topology approach to classify the critical points. Drawing inspiration from cq4 , which implies the allocation of topological charge to these critical points, we denote the temperature as $T=T\left(S,P,x^{i}\right)$. Based on this premise, the determination of the critical point of the black hole can be expressed as
$\left(\partial_{S}T\right)_{P,x^{i}}=0,\quad\left(\partial_{S,S}T\right)_{P,x^{i}}=0.$ (20)
The elimination of thermodynamic pressure leads to the derivation of a novel temperature function from equation (20), the next step entails the construction of a Duan’s potential cq4
$\Psi=\frac{1}{\sin\theta}T\left(S,x^{i}\right),$ (21)
here, the presence of an additional factor $1/\sin\theta$ contributes to the topological analysis. The application of Duan’s $\phi$-mapping theory is employed to introduce a new vector field denoted as $\phi=(\phi^{S},\phi^{\theta})$, where
$\phi^{S}=(\partial_{S}\Psi)_{\theta,x_{i}},\phi^{\theta}=(\partial_{\theta}\Psi)_{S,x_{i}}.$ (22)
Inspired by cq5 , in the case where $\theta$ equals $\frac{\pi}{2}$, the $\phi$ component of the vector field remains consistently at zero, allowing for a description of the topological flow as follows:
$J^{\mu}=\frac{1}{2\pi}\epsilon^{\mu\nu\lambda}\epsilon_{ab}\partial_{\nu}n^{a}\partial_{\lambda}n^{b},$ (23)
where $\partial_{\nu}=\frac{\partial}{\partial x^{\nu}}$ and $x^{\nu}=(t,r,\theta)$. The unit vector $n$ is represented by $n=(n^{1},n^{2})$, where $n^{1}=\frac{\phi^{S}}{\left\|\phi\right\|}$, $n^{2}=\frac{\phi^{\theta}}{\left\|\phi\right\|}$. The fulfillment of $\partial_{\mu}J^{\mu}=0$ is a prerequisite for the topological current. In addition, the topological charge linked to a parameter region $\sum$ can be expressed as follows:
$Q_{t}=\int_{\sum}j^{0}d^{2}x=\sum_{i=1}^{N}w_{i},$ (24)
the winding number $w_{i}$ determines the topological charge $Q$ at zero points $(\phi^{a}(x^{i})=0)$, with negative or positive values corresponding to a winding number of $-1$ or $+1$ respectively. The summation of charges at all critical points within a thermodynamic system defines the total topological charge $Q_{t}$, this fundamental concept holds great significance in the analysis of thermodynamic systems as it aids in comprehending their global characteristics and facilitating classification. By aggregating the charge associated with each critical point, we can derive a quantitative indicator to depict the system’s topology structure. The classification of distinct thermodynamic systems is determined based on the total topological charge. The aforementioned approach will be employed to investigate dyonic AdS black holes with QTE in the context of EGB gravity.
Figure 1: $d=4$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The green dot are $(r,\theta)=(2.449490,\pi/2)$, which represent the critical points $CP_{1}$. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve).
Figure 2: $d=5$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The blue dot are $(r,\theta)=(2.658528,\pi/2)$, which represent the critical points $CP_{1}$. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (blue curve), $C_{2}$ (brown curve).
Figure 3: $d=6,\alpha=0.1$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The green dot are $(r,\theta)=(1.711584,\pi/2)$, which represent the critical points $CP_{1}$. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve).
III Topology of critical points
In this section, we will investigate the categorization of critical points for black holes in all dimensions of space-time. By analyzing the metric of black holes in equation (5), we can deduce the thermodynamic function representing temperature without pressure
Figure 4: $d=6(\alpha=0.5,\beta=0.1,Q_{e}=1,Q_{m}=2)$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The colored dots are $(r,\theta)=(1.489048,\pi/2)$, $(r,\theta)=(2.706456,\pi/2)$ and $(r,\theta)=(1.873569,\pi/2)$ which represent the critical points $CP_{1}$, $CP_{2}$ and $CP_{3}$, respectively. Down panel:The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve) and $C_{3}$ (blue curve) and $C_{4}$ (black curve).
Figure 5: $d=6(\alpha=1,\beta=0.1,Q_{e}=1,Q_{m}=2)$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The colored dots are $(r,\theta)=(1.008427,\pi/2)$, $(r,\theta)=(3.609964,\pi/2)$ and $(r,\theta)=(3.284875,\pi/2)$ which represent the critical points $CP_{1}$, $CP_{2}$ and $CP_{3}$, respectively. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve) and $C_{3}$ (blue curve).
Figure 6: $d=7$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The blue dot are $(r,\theta)=(0.929668,\pi/2)$, which represent the critical points $CP_{1}$. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (blue curve), $C_{2}$ (green curve).
Table 1: Critical values at $Q_{e}=1,Q_{m}=2,\beta=0.1$ for dyonic AdS black holes with QTE in the EGB background.
$CP_{1}$ $CP_{2}$ $CP_{3}$
$r_{c}$ 2.449490 - -
($\alpha=0.5,d=4$) $T_{c}$ 0.043317 - -
$P_{c}$ 0.003316 - -
$r_{c}$ 2.658528 - -
($\alpha=0.5,d=5$) $T_{c}$ 0.063461 - -
$P_{c}$ 0.006193 - -
$r_{c}$ 1.711584 - -
($\alpha=0.1,d=6$) $T_{c}$ 0.138928 - -
$P_{c}$ 0.027903 - -
$r_{c}$ 1.489048 2.706456 1.873569
($\alpha=0.5,d=6$) $T_{c}$ 0.064736 0.064872 0.064642
$P_{c}$ 0.006536 0.006576 0.006353
$r_{c}$ 1.008427 3.609964 3.284875
($\alpha=1,d=6$) $T_{c}$ 0.055925 0.045933 0.045928
$P_{c}$ 0.050611 0.003311 0.003308
$r_{c}$ 0.929668 - -
($\alpha=0.5,d=7$) $T_{c}$ 0.113123 - -
$P_{c}$ 0.288091 - -
$\Psi$ $=\frac{1}{4\pi\sin\theta r_{+}\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)}$ (25)
×$\times\left[2(3-d)\left(2(d-5)(d-4)\alpha+r_{+}^{2}\right)\right.$
$\left.+r_{+}^{8-2d}\left(Q_{m}^{2}+\frac{r_{+}^{4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{2}}\right)\right].$
The vector field $\phi$ can be resolved into its components $(\phi^{r_{+}}$ and $\phi^{{\theta}})$, through appropriate manipulation, we can derive the outcome represented by
$\phi^{\theta}$ $=\frac{\cot\theta\csc\theta}{4\pi r_{+}\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)}\left[2(3-d)\right.$ (26)
×$\left.\times 2(d-5)(d-4)\alpha+r_{+}^{2}\right)+r_{+}^{8-2d}$
×$\left.\left.\times\left(Q_{m}^{2}+\frac{r_{+}^{4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{2}}\right)\right)\right],$
$\phi^{r_{+}}=\frac{\csc\theta}{4\pi\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)^{2}}[4(d-3)\alpha$ (27)
$\left.+r_{+}^{2}\right)+2r_{+}^{8-2d}\left(\frac{r_{+}^{4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{2}}\right.$
$\left.+Q_{m}^{2}\right)-2\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)$
×$\times\left(6-2d-\frac{(d-4)r_{+}^{6+4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{3}}\right.$
$+r_{+}^{6-2d}Q_{m}^{2}\left(+\frac{8d\beta\Gamma(d-1)^{2}r_{+}^{4+4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{3}}\right.$
$+4-d))+\frac{1}{r_{+}^{2}}\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)$
×$\times\left(-2(d-3)\left(2(d-5)(d-4)\alpha+r_{+}^{2}\right)\right.$
$+\left.\left.r_{+}^{8-2d}\left(Q_{m}^{2}+\frac{r_{+}^{4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{2}}\right)\right)\right].$
The normalized vector field is defined by
$n=(\frac{\phi^{r}}{\left\|\phi\right\|},\frac{\phi^{\theta}}{\left\|\phi\right\|}).$ (28)
Next, we investigate the topological charge linked to the critical point across various dimensions. Through analysis and examination of the topological structure, the topological charge $Q$ is assigned a specific value for each critical point, when encircling a single critical point $CP$ with a contour $C$, it is observed that its non-zero topological charge $Q$ signifies significant information associated with the enclosed area. Conversely, if contour $C$ lies outside of critical point, its topological charge becomes zero cq6 ; cq7 . If there are many critical points within the contour, the total topological charge $Q$ is obtained by aggregating the individual topological charges associated with each critical point. In the orthogonal plane, we employ the method of Duan’s $\varphi$-mapping theory to ensure that the constructed contour exhibits a positive value. We can transform the critical point on the $(S,\theta)$ plane to the $(r,\theta)$ plane using reparameterization. In this case, the critical point is situated at $(r_{0},\frac{\pi}{2})$
$\left\{\begin{matrix}r=a\cos\vartheta+r_{0},\\ \theta=b\sin\vartheta+\frac{\pi}{2}.\end{matrix}\right.$ (29)
where $\vartheta\subset(0,2\pi)$. After this transformation, we can express the deflection angle
$\Omega(\vartheta)=\int_{0}^{\vartheta}\epsilon_{ab}n^{a}\partial_{\vartheta}n^{b}d\vartheta.$ (30)
Naturally, the topological charge can be determined by the deflection angle as follows:
$Q=\frac{\Omega(2\pi)}{2\pi}.$ (31)
Now, let us investigate the classification of critical points pertaining to AdS black holes. Research has indicated that black holes demonstrate characteristic swallowtail behavior across various dimensions of space-time and undergo small/large BH phase transitions resembling those observed in van der Waals liquid/gas systems cq3 . Interestingly, in the case of $d=6$, when the coupling constant $\alpha$ falls within a particular parameter range $(\alpha\geq 0.5)$, a noteworthy phenomenon occurs where the black hole experiences a phase transition characterized by small/intermediate/large BH phases. Additionally, this leads to the emergence of a triple points naturally, signifying that the black hole possesses three critical points. In the case where the coupling constant $\alpha=1$, it can be observed that despite the presence of three critical points, there are indications of the triple point phase transition within the system as suggested by the existence of two swallowtails. We aim to topologically classify the critical points for the dyonic AdS BHs with QTE in the EGB background. Based on the data presented in TABLE I, it is clear that when $d$ takes values of 4, 5, 6 and 7, there exists only one critical point for the BH, which undergoes a small/large phase transition in the system. For the case of $d=4$, the characteristics of $\theta$ and radius $r_{+}$ are presented in relation to the vector field $n$ in FIG. 1. Identify a critical point located at $(\theta,r_{+})$=$(\frac{\pi}{2},2.449490)$. Create two contours, namely $C_{1}$ and $C_{2}$, where $C_{1}$ should enclose critical point, while no critical points should be enclosed by $C_{2}$ (refer to TABLE II for contour parameters). In addition, the values of $\Omega(\vartheta)$ are observed to be $-2\pi$ and 00, respectively. The topological charge at the critical point $CP_{1}$ is $-1$. In contrast, contour $C_{2}$ did not enclose any critical point, resulting in a topological charge of zero. Inspired by cq4 , $CP_{1}$ is classified as a conventional critical point, and the total charge is $Q=Q_{CP_{1}}=-1$. The presence of only one critical point in the system, which is called as a physical critical point. Similarly, in the case of space-time dimensions $d=5$ and $7$, we observe analogous outcomes as for $d=4$. The distinction lies in the fact that the coordinates of the physical critical points differ $(d=5,(\theta,r_{+})=(\frac{\pi}{2},2.658528),d=7,(\theta,r_{+})=(\frac{\pi}{2},0.929668))$. Further details can be found in FIG.2 and FIG.6.
Table 2: Contours at $Q_{e}=1,Q_{m}=2,\beta=0.1$ for dyonic AdS black holes with QTE in the EGB background.
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$
$a$ 0.07 0.07 - -
($\alpha=0.5,d=4$) $b$ 0.4 0.4 - -
$r_{0}$ 2.449490 3.4 - -
$a$ 0.07 0.07 - -
($\alpha=0.5,d=5$) $b$ 0.4 0.4 - -
$r_{0}$ 2.658528 3.8 - -
$a$ 0.15 0.15 - -
($\alpha=0.1,d=6$) $b$ 0.4 0.4 - -
$r_{0}$ 2.57753 2.4 - -
$a$ 0.15 0.15 0.15 1.1
($\alpha=0.5,d=6$) $b$ 0.4 0.4 0.4 0.5
$r_{0}$ 1.489048 2.706456 1.873569 2.09775
$a$ 0.15 0.15 0.15 -
6($\alpha=1,d=6$) $b$ 0.4 0.4 0.4 -
$r_{0}$ 1.008427 3.609964 3.284875 -
$a$ 0.15 0.15 - -
6($\alpha=0.5,d=7$) $b$ 0.4 0.4 - -
$r_{0}$ 0.929668 2 - -
Next, let’s analyze the situation where $d=6.$
$(i)$: $\alpha=0.1,\beta=0.1$
The presence of a conventional critical point for a black hole can be observed at coordinates $(1.711584,\pi/2)$, with a topological charge of $(Q=Q_{CP_{1}}=-1)$ in FIG.3. This result exhibits remarkable similarities to findings in other dimensions $(d=4,5,7)$.
$(ii)$: $\alpha=0.5,\beta=0.1$
[…]
$(iii)$: $\alpha=1,\beta=0.1$
We find that there are two swallowtails curves in the G-T diagram cq3 in the range of pressure $CP_{3}<P<CP_{2}$, we will discuss why the system has three critical points and only small/large BH phase transition occurs. The first observation is that the contours $C_{1}$ ,$C_{3}$ , and $C_{2}$ enclose the critical points $CP_{1}$ $(1.008427,\pi/2)$,$CP_{3}$ $(3.284875,\pi/2)$, $CP_{2}$ $(3.609964,\pi/2)$ respectively within the vector field in FIG.5. And the contours $C_{1}$ , $C_{2}$, and $C_{3}$ corresponding deflection angle $\Omega(\vartheta)=-2\pi,-2\pi$ and $2\pi$ for $\vartheta=2\pi$, respectively, this means that the black holes associated topological charge of $Q_{CP_{1}}=Q_{CP_{2}}=-$1, and $Q_{CP_{3}}=1$. As a result, the total topological charge can be determined by $Q=Q_{CP_{1}}+Q_{CP_{2}}+Q_{CP_{3}}=-1$. Naturally, the conventional critical point is denoted as $CP_{1}$ and $CP_{2}$, whereas the novel critical point is represented by $CP_{3}$. As the system undergoes small/large BH phase transition, it can be inferred that only $CP_{1}$ represents the conventional (physical) critical point. Furthermore, $CP_{2}$ and $CP_{3}$ are associated with non-physical critical points. In addition, the findings indicate that the critical point with topological charge of $-1$ can be considered as an indicator of the phase transition, whereas the novel critical point does not possess the ability to minimize Gibbs free energy and thus is defined as the critical point associated with non-physical characteristics ni1 . However, the conventional critical point $CP_{2}$ here lacks the ability to minimize the Gibbs free energy and is defined as a non-physical critical point. This suggests that the aforementioned conclusion cannot be universally applied to all systems.
Figure 7: a isobaric curves (colored solid curves) for dyonic AdS black holes with QTE in the EGB background. shown in the $T-r_{+}$ plane. The black curve is for the extremal points of the temperature. Up panel:$(\alpha=0.1)$,Middle panel:$(\alpha=0.5)$, Down panel:$(\alpha=1)$
IV Nature of Critical points
In this section, we will investigate the nature of the critical points, it can be observed that the black hole experiences small/large phase transition with only one critical point $(d=6,\alpha=0.1)$. Similar conclusions can be drawn for $d=4$, $5$, and $7$. Hence, we will primarily focus on examining the critical point of the black hole at $d=6$. To this end, the equation of state (12) is employed to generate the isobaric curves(represented by colored curve) and the critical temperature (depicted as a black curve) in FIG.7. We can find some interesting results as follows:
$(i)$ $\alpha=0.1,\beta=0.1$
Up panel: Each isopiestic curve corresponds to distinct maximum extreme points of temperature, the red curve representing $P=0.02$, the green curve represents $P=0.025$, the blue curve means $P_{CP1}=0.027903$, and the orange curve indicates $P=0.03$. Additionally, it is noteworthy that the conventional critical point $(Q_{CP1}=-1)$ coincides with the maximum temperature extreme point. We present a $T-S$ equal-area diagram in FIG.8, here $T_{p}$ denotes the temperature at which the black hole experiences a phase transition.
Figure 8: $T-S$ diagram of the Maxwell ’s equal area law $(P=0.025,T_{p}=0.1349,\alpha=0.1,d=6)$.
The application of Maxwell’s equal area law, we can eliminate the unstable region near $CP_{1}$ (referred to as the intermediate black hole branch). Moreover, our findings reveal that the isobaric curves are divided into distinct regions with varying phase numbers by the critical temperature, and the arrangement of phases in proximity to the critical point is determined based on these regions. Interestingly, an increase in pressure $(red\rightarrow orange)$ results in a reduction of isopiestic lines near $CP_{1}$. Consequently, $CP_{1}$ can be regarded as a phase annihilation point.
$(ii)$ $\alpha=0.5,\beta=0.1$
Middle panel: We can find that the isobaric curves $P_{CP3}=0.06353(purple)$; $P=0.006469(orange)$; $P_{CP1}=0.006536(blue)$; $P_{CP2}=0.0065757(red)$ exhibits temperature extreme points. We have observed that the maximum extreme temperature is associated with the green dots, which represent conventional critical points $CP_{1}$ and $CP_{2}$. On the other hand, the blue point (novel critical point $CP_{3}$) accurately represents the minimum extreme temperature. We can note that the region of $T-S$ is depicted in FIG.9, with $T_{p1}$ and $T_{p2}$ denoting the temperature of the black hole during small to intermediate and intermediate to large phase transitions, respectively. By applying Maxwell’s equal area law, we can effectively eliminate unstable regions near $CP_{1}$ and $CP_{2}$; however, this approach may not be successful for unstable regions near $CP_{3}$. These findings align perfectly with the results cq4 . Furthermore, as pressure increases (indicated by orange-red), we notice a decrease in the number of regions surrounding $CP_{1}$ and $CP_{2}$, as well as a decrease in the number of regions near $CP_{3}$.
Figure 9: $T-S$ diagram of the Maxwell ’s equal area law $(P=0.006492,T_{p1}=0.064718,T_{p2}=0.0647425,\alpha=0.5,d=6)$.
Figure 10: $T-S$ diagram of the Maxwell ’s equal area law $(P=0.025,T_{p}=0.05328,\alpha=1,d=6)$.
We designate $CP_{3}$ as the phase generation point while referring to critical points $CP_{1}$ and $CP_{2}$ as phase annihilation points.
$(iii)$ $\alpha=1,\beta=0.1$
Down panel: We can find that the isobaric curves $(P_{CP3}$=$0.03308(purple)$; $P$=$0.003309(orange)$; $P_{CP2}$=$0.003311(blue)$; $P=0.025(green);(P_{CP1}=0.050611(red))$ exhibit temperature extreme points. Similarly, the green dots ($CP_{1}$ and $CP_{2}$) correspond to the point of maximum extreme temperature point, while the blue dot (novel critical point $CP_{3}$) accurately corresponds to the point of minimum extreme temperature point. Based on research findings cq3 , it has been observed that there is no occurrence of a first-order phase transition in close proximity to the conventional critical point $CP_{2}$, which can also be found in we2 ; we7 . In FIG.10, we give the temperature $(T_{p}=0.05328)$ at which the $P=0.025$ black hole undergoes a phase transition. Notably, the application of Maxwell’s law of equal area can effectively remove the unstable region near $CP_{2}$. However, it is not suitable for indicating a first-order phase transition. Therefore, an alternative classification should be employed, it can be inferred that $CP_{2}$ may serve as a phase annihilation point.
V Conclusions
In summary, we employ the Duan ’s $\varphi$-mapping theory to examine the topological structure of thermodynamics in black holes. The categorization of critical points for black holes is carried out across diverse dimensions of space-time. In the scenario where $d=6$, the black hole will have triple points when $\alpha$ specific value is assumed by the GB coupling constant. Our focus lies on assessing how this coupling constant impacts the critical point. This investigation contributes to a more profound comprehension of EGB modified black holes. The primary highlights can be summarized as follows:
$(i)$ Black holes have a critical point in any space-time dimension, the presence of a small/large BH phase transition gives rise to a conventional (physical) critical point, attributed to the total topological charge $-1$. Different $d$ will lead to different positions of the critical point in vector field $n$. Initially, as d increases, the critical point also increases before eventually decreasing again in TABLE I.
$(ii)$ For $d=6$, we will find that when the coupling constant satisfies $\alpha=0.5$ and $1$, the system will appear triple points, but the phase structure is completely different. More specifically, the three critical points can be classified into two conventional (physical) critical points, $(CP_{1}$ and $CP_{2})$, as well as a novel critical point $(CP_{3})$, the novel critical point is not a physical point because it does not minimize the Gibbs free energy. When a = 1, the system has only one traditional ( physical ) critical point $(CP_{1})$. In both cases, the total topological charge is $-1$.
$(iii)$ From the isobaric curve and critical temperature in FIG.7, we can observe that the conventional critical point is located at the maximum extreme point of temperature, and the unstable region in close proximity to it can be eliminated through the application of Maxwell’s equal area law. It is found that the critical point $CP_{3}$ is no longer applicable. Noted that the novel critical point appears at the minimum extreme point of temperature. The results also reveal a decrease in the number of phases along the isobaric curve at the conventional critical point as pressure increases $(\alpha=1)$, while the number of phases increases at the novel critical point. Therefore, we define the $CP_{2}$ as the phase annihilation point $(\alpha=1)$.
Acknowledgements.
This work is supported by the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07, QJJ-[2022]-314), and the National Natural Science Foundation of China (Grant Nos. 12265007 and 12264061).Bekir Can Lütfüoğlu is grateful to Excellence project PřF UHK 2211/2023-2024 for the financial support.